March 19th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • setting up and testing cable attenuation with LED
      • Chuck grounded the HCAL frame ( not causing noise ) and

racks grounded

      • completed GUI for connection for HCAL - will copy it to enpcamsonne
      • Bill will modify more discriminator for threshold
      • ask electronics group if they can make adapter for DAQ
      • Juan Carlos wants to know if we have a calibration of DAC input to discriminator threshold - need to find out -

Brad 10 to 1 for previous version - and offset was added - will need to check

      • talked with Ed about F1 v1, can use the clock from TI but need different configuration file. Program exists for v2 and v3, need to adapt for v1 and plug clock from TI to F1 - HCAL would like 0.25 ns timing resolution
    • BB Shower
      • making cables for triggering monitoring system need 17
      • debugging some channels, preshower ? left side seems to have additionnal pedestal noise
      • issue with Fastbus : need to take snapshot of errors and post snippet of bad data - trigger rate 100 to 250 Hz - recording event - no errors - all 0 in raw file - goes away after power cycling - not sure when it started to happen -

happened in between taking runs - maybe 15 to 20 mins - will send first run number when that happened to Bryan

      • DSG made 200 cables for FADC ready
      • 12 FADCs from Chris should be available
    • BB Hodo
      • data for entire detector - all hardware issues
      • analysis scripts being debugged - will use for gain matching
      • HV CAEN : EPICS - Hall B software can talk to it, use EPICS and Hall C style tk gui -
      • looking at TDC for scaler display - can adapt for other detectors with F1
    • GRINCH
      • software issue installing SBS-offline, issue with analyzer. Everything set up now
      • Gas will be delivered today
      • concerns using V792 - will use Fastbus 1881 -
    • LHRS
      • Maria and Bob continue working
      • Issue with FADC readout in mode 10 - segmentation fault with waveform - comparing with Hall - /proc/meminfo can give memory info - top should work - could compare with Bigbite FADC readout list ?
      • looked at Bigbite DAQ with Arun
      • Need to switch to HCal FADC firmware ?
    • CDet
      • sbs2 kept rebooting - managed to boot but plan to move to spare computer
      • issue running startcoda on sbs2 when it can boot
      • 5 out 6 modules cable up - started taking data
    • GEM INFN
      • no progress - still need to set back up
    • GEM TEDF
      • INFN : trying to setup pulser trigger with SSP readout
      • trying to run UVA and INFN GEM together :

8 MPDs for INFN - added each MPD one at a time - was working taking data - started added MPD for UVA - but SSP timeout issue - but issue still there when MPDs UVA removed might be issue with many MPDs initialization - looking with VME code : still see MPD initialization issue so might not SSP issue - data with pulser data and with HV but DAQ unstable. Ben will come today to check if there is a voltage drop at initialization

      • LV supply on network : can control with web interface - need pyvisa installed so can control the supply
    • GEM UVA
      • still have APV baseline fluctuations - tested Nilanga suggestion : use single LV for each plane - still see same behavior - might come from aluminum frame : looks better with separate chamber - still investigating
  • DAQ bunker
    • Issue with crane in Hall A : weldment will not be moved soon - Jack and Jessie agree that we can move individual relay racks - Bigbite old weldment in ESB